OpenStudy (anonymous):

solve this system of equations : show steps x^2+y^2=36 y=x+2

myininaya (myininaya):

so y=x+2 plug this into the first equation to get \[x^2+(x+2)^2=36\]

myininaya (myininaya):

solve this for x

myininaya (myininaya):

\[x^2+x^2+4x+4=36\]

OpenStudy (anonymous):

it is a circle and a line

myininaya (myininaya):

\[2x^2+4x+4-36=0\] There that should help

OpenStudy (anonymous):

oh..okay :D.. i got the first part, was jus confused on the second..lols thnks!!!